OpenStudy (anonymous):

need to combine like terms and simplify this problem 1+3/4(10t-3)+5/8(t+1/10)

Directrix (directrix):

Would you click on the blue DRAW button below and write the equation by hand? It is hard to tell which terms are in the denominators and which are in the numerators. Thanks.

OpenStudy (anonymous):

\[1+\frac{ 3 }{ 4 }(10t-3)+\frac{ 5 }{ 8 }(t+\frac{ 1 }{ 2 }\]

Directrix (directrix):

>5/8(t+1/10) from the problem posted does not match with the typed problem. Should the problem end in 1/2 or 1/10 ?

OpenStudy (anonymous):

show i rewrite it

OpenStudy (anonymous):

\[1+\frac{ 3 }{ 4 }(10t-3)+\frac{ 5 }{ 8 }(t+\frac{ 1 }{ 10 })\]

Directrix (directrix):

I would use the Distributive Property 1 + (3/4) * (10t - 3) + (5/8)* (t+(1/10)) = 1 + (3/4) * (10t) + (3/4) * (-3) + (5/8)* (t) + (5/8) * (1/10)

OpenStudy (anonymous):

ok then what

Directrix (directrix):

I would do each calculation by parts: (3/4) * (10t) = (30t)/4 (3/4) * (-3) = (-9)/4 (5/8)* (t) = ? (5/8) * (1/10) = ? ------------ After that, we have to add those together and that sum is added to the 1.

OpenStudy (aonz):